Trey has two dependents, his daughters, ages 14 and 17, at year-end. Trey files a joint return with his wife.

What amount of child credit will Trey be able to claim for his daughters under each of the following alternative situations? Use Exhibit 8-8.

c. His AGI is $420,100, and his daughters are ages 10 and 12.
    

EXHIBIT 8-8 Child Tax Credit Phase-Out Threshold

Filing Status

Phase-Out Threshold

Married filing jointly

$400,000

Married filing separately

200,000

Head of household and single

200,000
Answer #1

The filing status of Trey is married filing jointly. Applicable threshold would be $ 400,000. Since, his AGI is $ 420,000 that is above the threshold, he is eligible for child credit with certain reductions for amount above phasout.

Child tax credit is applicable to eligible child. Assuming all the other conditions is complied, the eligible child is under the age of 17 i.e. 16 or younger. Hence, both child is eligible for child tax credit.

Hence. before phase-out, Trey has a $2,000 child tax credit for each of his qualifying children ages 10 and 12 (under 17) at year end, total tax credit being $ 4,000. The tax credit completely phases out at $ 440,000

Now, computing the child tax credit using the steps below

AGI above threshold = 420,100 – 420,000 = $ 20,100

This has to be divided by $1000 = 20,100/1000 = 21.

After phase out, every $ 50 dollar shall be reduced for every $ 1000 increase in AGI above phaseout.

Total decline in tax credit = 50 X 21 = $ 1050

Hence, child tax credit after reduction = 4000 – 1050 = $ 2,950.